Operations Report – December 1st

SOL: 4

Name of person filing report: Ashley Kowalski

Non-nominal systems: N/A

Notes on non-nominal systems: N/A


Spirit rover used: Yes
Hours: 201.8 (Start), 201.9 (End)
Beginning charge: 100%
Ending charge: 92%
Currently charging: No

Opportunity rover used: No
Hours: N/A
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Curiosity rover used: No
Hours: N/A
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Perseverance rover used: Yes
Hours: 250.3 (Start), 250.3 (End)
Beginning charge: 100%
Ending charge: 90%
Currently charging: No

General notes on rovers: No additional notes/comments.

Summary of Hab operations:
WATER USE: 30 gallons used today, 110 gallons used in total thus far
Water (static tank): 440 gallons
Static tank pipe heater (on or off): On
Static tank heater (On or off): On
Toilet tank emptied: No.

Summary of internet: No issues with the internet. Connection is good.

Summary of suits and radios:

The Crew Engineer (Ashley Kowalski) and GreenHab Officer (Matthew Eby) performed repairs on EVA suit #7. Ashley and Matt located the soldering iron, solder, and wire in the RAM and performed the repairs in that module. In order to remove the old solder, they used the soldering iron to melt it off. This enabled them to push new wires through the connection openings and then resolder new wires to that area. They were able to manually peel off any of the additional glue around the previous connections. After cleaning up the connection points, all wire connections were successfully reconnected to the old charging socket (we used EVA suit #8 as an example to follow). Please see attached photo displaying the repaired connections (please see photo 001 attached). However, even with all of the wire connections reconnected correctly, the fans were still not turning on. Therefore, Ashley and Matt tested all of the connections with the multimeter and noticed that even though the battery was reading ~13 V, the battery power was not transferring to the on/off switch. Upon further investigation, they realized that the fuse was blown (please see photo 002 attached). Thus, they located spare fuse parts in the EVA suit repair kit and put a new fuse in the system, which solved the problem. The fans turned on, the charging port was reconnected, and EVA suit #7 is now cleared for use.

One additional point that was noticed during the repair was that the lower connection on the on/off switch (the one that is circled in photo 003 attached; it is the one that connects the battery power to the fans) is a bit loose. Ashley and Matt attempted to tighten this connection, but it is still a bit loose. Why is this important? They believe that the looseness of this connection might have been the reason that the fuse blew. As you can see in the video uploaded to the Google drive link, when that connection is in one position (not pushed all the way down, but pushed slightly inward), it reads an incorrect value on the multimeter. When the connection IS pushed all the way down (pushed outward), the correct, expected value is read on the multimeter. Ashley and Matt recommend replacing this switch in the future as this connection may cause problems for future crews.

Finally, the two walkie talkies mentioned previous reports (one from the upstairs set and one from the downstairs set) are still not working.

Summary of GreenHab operations: Today, the crew joined the GreenHab Officer in the greenhouse and helped with planting carrots and onions. Additional information can be found in the GreenHab Report.
WATER USE: ~23 gallons (~5 gallons today)
Heater: On
Supplemental light: Off
Harvest: N/A

Summary of ScienceDome operations: N/A
Dual split: No heat or AC used. All systems off.

Summary of RAM operations: A soldering iron, solder, wire, and a spare EVA suit fuse were used in the RAM to repair EVA suit #7. The heater was turned on while the Crew Engineer and GreenHab Officer were in the RAM performing the repairs. The heater was turned off and unplugged upon exiting the RAM.

Summary of any observatory issues: No observatory issues.

Summary of health and safety issues: No health and safety issues.

Questions, concerns and requests to Mission Support: The Crew Engineer provided Mission Support with the dimensions of the microwave grease filter. Also, when the wind picked up this afternoon before the crew was able to attach the additional security rope to the front airlock outer door, the door swung upon. The crew received permission to re-secure the door. Finally, we received the air filter for the furnace, but the Crew Engineer has not yet installed it.

Copyright © The Mars Society. All rights reserved. | Main Site